OpenStudy (anonymous):

Write an equation in point-slope form of the line through point J(–5, 6) with slope –4. y - 6=-4(x-5) y - 6=4(x-5) y + 6=-4(x-5) y - 6=-4(x+5)

OpenStudy (anonymous):

y - 6=-4(x+5)

OpenStudy (anonymous):

y + 6=-4(x-5)

OpenStudy (anonymous):

so witch one is it?

OpenStudy (anonymous):

y - 6=-4(x+5)

OpenStudy (anonymous):

y - 6=-4(x+5)

OpenStudy (anonymous):

yay hero :D thx everybody